# Framing Political Bias in Multilingual LLMs Across Pakistani Languages

### v2

@@ 5.2 Framing Bias Analysis
Stylistic Bias: Lexical polarity analysis highlights how LLMs stylistically frame within Pakistani political discourse. Figure 14 shows some of society’s most sensitive and debated issues. Same-sex marriage rights stand out with the most polarized coverage—yet interestingly, the sentiment leans slightly positive, hinting at a complex and emotionally charged discourse. Language policy, welfare and charity, and education follow closely, marked by passionate debate but generally hopeful tone. On the other hand, deeply rooted religious and moral issues like abortion, blasphemy laws, and the death penalty show intense division and overwhelmingly negative sentiment. When it comes to government performance, coverage tends to be both critical and sharply divided—reflecting growing public frustration as shown in Figure 17. Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression. For more analysis and results see Appendix C.
@@ 6.1 Model-Level Bias Interpretation
Figure  8 highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini. The results highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini. To further interpret how political bias manifests in model behaviour, we conducted a detailed, multidimensional analysis of the DeepSeek model’s visual and quantitative breakdowns, which are provided in the Appendix Figure 21.
@@ 6.2 Cross-Cultural model insights
Our analysis reveals distinct model behaviors across cultural-linguistic contexts. GPT models show high cultural sensitivity, shifting from libertarian-left in English to authoritarian-left in Pakistani languages (Figure 4). Claude remains consistently libertarian with progressive sentiment. Open-source models vary—Deep Seek offers stable entity selection, while Gemini shows notable ideological shifts. These findings challenge universal bias mitigation, emphasizing that effective global AI requires balancing fairness with cultural nuance—something current models often fail to achieve.
@@ 7 Conclusion
This study presents the first large-scale investigation of political bias in large language models across five low-resource yet widely spoken Pakistani languages: Urdu, Punjabi, Sindhi, Pashto, and Balochi. We introduce a novel, culturally adapted evaluation framework that integrates ideological stance scoring with discourse-level framing analysis—capturing both what the model says and how it says it. Our findings uncover systematic, language-conditioned shifts in political bias, demonstrating how linguistic and cultural context significantly shapes model behavior. By bridging the Political Compass Test (PCT) with narrative framing dimensions, our methodology offers a reproducible and extensible approach for diagnosing political bias in multilingual LLMs. This work fills a critical gap in multilingual NLP and contributes a practical auditing tool for building culturally grounded, fair, and context-aware AI systems.
@@
This study has several methodological and scope-related limitations. Our analysis is restricted to five Pakistani languages, omitting other regional or global languages that may exhibit different ideological patterns. We also exclude high-resource languages like French or Arabic, which could have served as cultural counterfactuals to help distinguish between linguistic and cultural influences in LLM pretraining. While we adapt the Political Compass Test (PCT), its original design is rooted in Western political thought and may not fully capture culturally specific structures like Islamic jurisprudence or tribal governance. Additionally, our sentiment and entity analysis tools are primarily trained on English, potentially reducing accuracy and missing localized expressions. We do not explore prompt-induced framing bias, which may affect stance independently of model ideology. Finally, the static nature of training data limits our ability to assess real-time political shifts. Future work should explore culturally sensitive bias mitigation for low-resource settings.
@@ A.1 Dataset Contribution
Our research provides a valuable multilingual dataset that spans five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies. For the Political Compass Test, the statement are translated into five languages shown in the Table 28. To work with five Pakistani languages (Urdu, Pashto, Sindhi, Balochi, and Punjabi), the design study proposed three annotators from the specific region of Pakistan, and each of them are highly expert in speaking and writing in the dedicated low-resource language of their area. The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. We chose the Method: Triple-verified by native speakers. All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions. The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  9. The generated response on 62 culturally adapted political statements in all five languages as shown in Table 28, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 3. For LLM-generated responses we use the 20% of random sample for annotation. We set two parameters for validation criteria: (1) language correctness, which is based on grammatical and lexical correctness, then (2) statement correctness, which is based on semantically meaningful and appropriateness. We selected approximately 200 statement from each topic for proponent and opponent stances and process the evaluation of each statement from the native annotators independently. For each statement, annotators will make two judgments:
@@ A.3 Translation Assignment Procedure
[…]
@@
Table 1: Summary Statistics for Political Compass Test Translations and Headline Generation Corpus
@@ Appendix B Technical details
B.1 Stance Detection vs. Our Ideological Mapping
@@
B.2 Prompting Strategy and Reproducibility
We adopt a zero-shot prompting strategy for all LLM generations to avoid bias from exemplars. For each of the 62 PCT statements, models are asked to respond in the specified language using a Likert-style agreement format. For framing analysis, we generate 1,000 headlines per stance–topic pair using bilingual prompts that explicitly request either supportive or opposing tone (see Appendix 21). Frame classification is conducted using GPT-3.5-turbo with a fixed schema based on Boydstun’s taxonomy. Named entity and sentiment analyses are automated via mDeBERTa and XLM-RoBERTa. All code and prompts are shared publicly for reproducibility.
B.3 Model Details
Table 2 provides an overview of the language models used in our bias evaluation. It includes both closed-source and open-source models, detailing their type, estimated parameter sizes, and architectures. Hyperlinked model names direct to official documentation or repositories, enabling transparency and reproducibility for further comparative analysis.
@@ B.3 Model Details
Table 2: Overview of Language Models Used in Bias Evaluation
@@
XLM-RoBERTa-large Open-source 550M Encoder
@@
B.4 On Bias and Limitations of mDeBERTa
@@
B.5 Stance Classification Details
@@ B.6 Standard PCT Automation
B.7 Language Focus and Fine-tuning Scope.
@@ C.2 Ideological Leanings of LLMs Across Political Topics
Figure 15: Scatter plot showing how model confidence relates to prediction uncertainty. When models are highly confident, their bootstrap error can remain significant. The trend line ($y=1.15x-0.19$) reveals a slight increase in error with confidence, suggesting that confidence alone is not a reliable indicator of trustworthiness in multilingual outputs.
@@ C.4 Boydstun Framing Model Confidence
Figure 15 illustrates the relationship between model confidence and the intensity of Boydstun Framing in responses across political discourse. Each point represents a model output, plotted by its confidence score (X-axis) and normalized framing intensity (Y-axis), with color indicating the dominance of a particular frame dimension (Frame Percentage). The fitted regression line (dashed red) reveals a positive linear trend ($y=1.15x-0.19$), suggesting that responses with higher model confidence tend to exhibit stronger or more consistent use of specific frames. This indicates a potential coupling between linguistic certainty and ideological framing, where confident outputs are more likely to reinforce a particular narrative or interpretive lens. The trend supports the hypothesis that framing is not incidental, but may become more pronounced when the model generates responses it deems more certain—highlighting the need for deeper scrutiny in high-confidence predictions when auditing bias in LLM outputs.
@@ C.7 Political Bias Insights Through Our Framework
The lowest diagrams provide insight on lexical polarity rates across topics, presenting a 35.00% positive rate for proponent stances versus -26.00% for opponent positions. Meanwhile, the right-side plotted graph displays the model’s stance on specific political issues within a coordinate system, with topics like education and religious minorities appearing in supportive positions, while others like blasphemy laws display opposition, as shown in Figure 11.
Finally, the flowchart illustrates how political stance detection leads to bias measurement through linguistic investigation, showing how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.
@@ C.9 Does Model Size Correlate with Political Neutrality?
Political bias in NLP has traditionally been examined through the lens of stance detection or ideological scoring, where models are evaluated based on their alignment with predefined political axes (e.g., liberal vs conservative, authoritarian vs libertarian). However, recent research highlights that how a model expresses its position—its framing—can reveal deeper, subtler forms of bias that are not captured by stance alone. Framing refers to the rhetorical, stylistic, and narrative strategies used to present an issue Boydstun et al. (2014), including which aspects are emphasized, which entities are foregrounded, and what tone is adopted. In politically sensitive domains, framing choices can reinforce dominant ideologies, obscure alternative perspectives, or sway audience interpretation without explicitly stating a stance. Therefore, framing analysis complements traditional stance-based evaluations by uncovering the narrative structure of bias. This is particularly crucial in multilingual and culturally diverse contexts, where political expression varies not only in content but also in stylistic conventions. Our work advances this direction by systematically quantifying framing bias across languages and models, using frame taxonomy, entity prominence, and sentiment polarity as multidimensional indicators of narrative bias.
C.11 Liberal-Leaning Tendencies in Political Topics.

### v3

@@ 5.2 Framing Bias Analysis
Stylistic Bias: Lexical polarity analysis highlights how LLMs stylistically frame within Pakistani political discourse. Figure 14 shows some of society’s most sensitive and debated issues. Same-sex marriage rights stand out with the most polarized coverage, yet interestingly, the sentiment leans slightly positive, hinting at a complex and emotionally charged discourse. Language policy, welfare and charity, and education follow closely, marked by passionate debate but generally hopeful tone. On the other hand, deeply rooted religious and moral issues like abortion, blasphemy laws, and the death penalty show intense division and overwhelmingly negative sentiment. When it comes to government performance, coverage tends to be both critical and sharply divided reflecting growing public frustration as shown in Figure 17. Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression. For more analysis and results see Appendix C.
@@ 6.1 Model-Level Bias Interpretation
Figure  7 highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini. The results highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini. To further interpret how political bias manifests in model behaviour, we conducted a detailed, multidimensional analysis of the DeepSeek model’s visual and quantitative breakdowns, which are provided in the Appendix Figure 21.
@@ 6.2 Cross-Cultural model insights
Our analysis reveals distinct model behaviors across cultural-linguistic contexts. GPT models show high cultural sensitivity, shifting from libertarian-left in English to authoritarian-left in Pakistani languages (Figure 4). Claude remains consistently libertarian with progressive sentiment. Open-source models vary Deep Seek offers stable entity selection, while Gemini shows notable ideological shifts. These findings challenge universal bias mitigation, emphasizing that effective global AI requires balancing fairness with cultural nuance something current models often fail to achieve.
@@ 7 Conclusion
This study presents the first large-scale investigation of political bias in large language models across five low-resource yet widely spoken Pakistani languages. We introduce a novel, culturally adapted evaluation framework that integrates ideological stance scoring with discourse-level framing analysis capturing both what the model says and how it says it. Our findings uncover systematic, language-conditioned shifts in political bias, demonstrating how linguistic and cultural context significantly shapes model behavior. By bridging the PCT with narrative framing dimensions, our methodology offers a reproducible and extensible approach for diagnosing political bias in multilingual LLMs. This work fills a critical gap in multilingual NLP and contributes a practical auditing tool for building culturally grounded, fair, and context-aware AI systems.
@@
This study has several methodological and scope-related limitations. Our analysis is restricted to five Pakistani languages, omitting other regional or global languages that may exhibit different ideological patterns. We also exclude high-resource languages like French or Arabic, which could have served as cultural counterfactuals to help distinguish between linguistic and cultural influences in LLM pretraining. While we adapt the Political Compass Test (PCT), its original design is rooted in Western political thought and may not fully capture culturally specific structures like Islamic jurisprudence or tribal governance. Additionally, our sentiment and entity analysis tools are primarily trained on English, potentially reducing accuracy and missing localized expressions. We do not explore prompt-induced framing bias, which may affect stance independently of model ideology. Finally, the static nature of training data limits our ability to assess real-time political shifts. Finally, while deterministic decoding ensures reproducibility, it limits robustness by providing only one output per prompt. Future work will address this by sampling multiple generations and applying statistical aggregation. Our approach is the reliance on GPT-3.5-turbo as the frame dimension classifier, which, despite human verification of a 20% sample confirming its reliability ($\kappa>0.7$), may still introduce subtle biases or misclassifications that future work should address with more diverse or human-supervised classifiers. Future work should explore culturally sensitive bias mitigation for low-resource settings.
@@ A.1 Dataset Contribution
Our research provides a valuable multilingual dataset that spans five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies. For the Political Compass Test, the statement are translated into five languages shown in the Table 28. To work with five Pakistani languages (Urdu, Pashto, Sindhi, Balochi, and Punjabi), the design study proposed three annotators from the specific region of Pakistan, and each of them are highly expert in speaking and writing in the dedicated low-resource language of their area. The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. We chose the Method: Triple-verified by native speakers.
All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions.
The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  8. The generated response on 62 culturally adapted political statements in all five languages as shown in Table 28, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 3. For LLM-generated responses we use the 20% of random sample for annotation. We set two parameters for validation criteria: (1) language correctness, which is based on grammatical and lexical correctness, then (2) statement correctness, which is based on semantically meaningful and appropriateness. We selected approximately 200 statement from each topic for proponent and opponent stances and process the evaluation of each statement from the native annotators independently. For each statement, annotators will make two judgments:
@@ A.3 Translation Assignment Procedure
$\kappa$ $<0:\text{Poor agreement}$
$0.01\leq\kappa$ $\leq 0.20:\text{Slight agreement}$
$0.21\leq\kappa$ $\leq 0.40:\text{Fair agreement}$
$0.41\leq\kappa$ $\leq 0.60:\text{Moderate agreement}$
$0.61\leq\kappa$ $\leq 0.80:\text{Substantial agreement}$
$0.81\leq\kappa$ $\leq 1.00:\text{Almost perfect agreement}$
@@
Table 1: Human Interannotator Correctness Summary Statistics for Political Compass Test Translations and Headline Generation Corpus
@@ Appendix B Technical details
B.1 Language and Political Identity in Pakistan
The relationship between language and political identity in Pakistan is firmly established in sociolinguistic scholarship and is consistent and validated by our findings. Classic work by Rahman (1996, 2011) documents how languages in Pakistan have served as enduring political symbols; Urdu functions as a marker of national and religious unity, while Sindhi, Punjabi, Pashto, and Balochi as anchors of ethnic and political resistance. Subsequent studies Umrani and Bughio (2020); Abbas and Bidin (2022) emphasize that language in Pakistan is not merely communicative but constitutive of political identity, shaping mobilization, rights claims, and policy contestation. In the study,Abbas and Bidin (2022) highlights that speakers of indigenous languages actively treat their mother tongues as identity markers, and that language policies have often been intermingled with covert political goals, reinforcing social stratification and contributing to historical conflicts. These sociolinguistic foundations motivate a culturally grounded bias evaluation: if languages are carriers of political identity, then multilingual probing should reveal systematic stance variation rather than translation artifacts. As we show empirically, Urdu responses tend to lean more liberal, while Pashto and Balochi skew conservative, patterns that align with documented orientations of their respective communities Rahman (2011); Umrani and Bughio (2020); Abbas and Bidin (2022).
B.2 Why adapt PCT for Pakistani languages?
The Political Compass Test (PCT) has been critiqued for its Western centrism and prompt sensitivity Röttger et al. (2024), yet it remains one of the few instruments systematically probing political orientation across ideological axes. Recent work has shown that, with careful cultural adaptation, PCT-style frameworks can yield meaningful insights in multilingual contexts, including Bangla Thapa et al. (2023), cross-regional evaluations Helwe et al. (2025), and large-scale comparative studies Bang et al. (2024).
Building on this line of research, we introduce the first culturally adapted PCT for five low-resource Pakistani languages. Our framework goes beyond stance mapping by integrating multi-level framing analysis capturing not only what positions LLMs adopt but also how they are rhetorically expressed through policy frames, entities, and lexical polarity. Validation rests on two pillars: (i) high inter-annotator agreement on translation and verification tasks, and (ii) systematic cross-language stance variation consistent with well-documented sociolinguistic patterns in Pakistan Rahman (2011); Umrani and Bughio (2020).
B.3 Stance Detection vs. Our Ideological Mapping
@@ B.2 Why adapt PCT for Pakistani languages?
B.4 Computational Resources:
This study required substantial computational resources to evaluate political bias across multilingual LLMs. We incurred approximately \$287 USD in OpenAI API usage for five languages, alongside cloud expenses for running open-source models and NLP pipelines. The total budget was around \$350 USD. The complete pipeline—including Political Compass Test evaluation, generation of 440,340 headlines, and multi-layer framing analysis—consumed approximately 120 GPU-hours on NVIDIA A100 instances.
Hyperparameter Settings: We ensured consistency across all model generations by using a zero-shot multilingual setup with fixed decoding parameters: temperature $T=0.0$ for deterministic outputs in bias-sensitive tasks and $T=0.5$ for controlled variation, with top-$p=1.0$ and a maximum token length of 150. For bias-sensitive evaluation, we adopt deterministic decoding ($T=0.0$) to eliminate randomness and ensure reproducibility. While this setting yields a single deterministic output per prompt, future extensions will incorporate multi-sample prompting and aggregation for robustness.
B.5 Model Details
The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o, Claude, and Gemini 1.5 Pro. The open-source models include: Mistral, DeepSeek, RoBERTa-large, RoBERTa-base, and BERT-large, model. Table 2 provides an overview of the language models used in our bias evaluation. It includes both closed-source and open-source models, detailing their type, estimated parameter sizes, and architectures. Hyperlinked model names direct to official documentation or repositories, enabling transparency and reproducibility for further comparative analysis.
@@ B.5 Model Details
BERT-base Open-source 110M Encoder
@@
XLM-RoBERTa-large Open-source 550M Encoder
Table 2: Overview of Language Models Used in Bias Evaluation
B.6 On Bias and Limitations of mDeBERTa
@@
B.7 Stance Classification Details
@@ B.8 Standard PCT Automation
B.9 Language Focus and Fine-tuning Scope.
@@ C.2 Ideological Leanings of LLMs Across Political Topics
Figure 15: Scatter plot showing how model confidence relates to prediction uncertainty. When models are highly confident, their bootstrap error can remain significant. The trend line (y=1.15x0.19y=1.15x-0.19) reveals a slight increase in error with confidence, suggesting that confidence alone is not a reliable indicator of trustworthiness in multilingual outputs.
@@ C.4 Boydstun Framing Model Confidence
Figure 15 illustrates the relationship between model confidence and the intensity of Boydstun Framing in responses across political discourse. Each point represents a model output, plotted by its confidence score (X-axis) and normalized framing intensity (Y-axis), with color indicating the dominance of a particular frame dimension (Frame Percentage). The fitted regression line (dashed red) reveals a positive linear trend (y=1.15x0.19y=1.15x-0.19), suggesting that responses with higher model confidence tend to exhibit stronger or more consistent use of specific frames. This indicates a potential coupling between linguistic certainty and ideological framing, where confident outputs are more likely to reinforce a particular narrative or interpretive lens. The trend supports the hypothesis that framing is not incidental, but may become more pronounced when the model generates responses it deems more certain—highlighting the need for deeper scrutiny in high-confidence predictions when auditing bias in LLM outputs.
@@ C.7 Political Bias Insights Through Our Framework
The lowest diagrams provide insight on lexical polarity rates across topics, presenting a 35.00% positive rate for proponent stances versus -26.00% for opponent positions. Meanwhile, the right-side plotted graph displays the model’s stance on specific political issues within a coordinate system, with topics like education and religious minorities appearing in supportive positions, while others like blasphemy laws display opposition, as shown in Figure 10. Finally, the flowchart illustrates how political stance detection leads to bias measurement through linguistic investigation, showing how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.
To ensure Prompt Reproducibility and Robustness our ideological measurements are not artifacts of a single instruction formulation, we conduct a systematic prompt reproducibility analysis. Each Political Compass Test (PCT) item is evaluated using five distinct prompt variants that preserve the same semantic task eliciting agreement or disagreement with a political statement, while varying instructional framing, contextual emphasis, and response constraints. Specifically, the prompt variants differ along three controlled dimensions: (i) reasoning style (opinion-based vs. analytical), (ii) contextual grounding (generic vs. Pakistan-specific), and (iii) response format constraints (free-form vs. fixed sentence length). All prompt templates are provided in Table 4.
Metric Analysis
PCT statements 62
Items with valid predictions across all prompts 60
Number of prompt variants 5
Mean per-item variance across prompts 0.068
Std. per-item variance 0.092
Cohen’s $\kappa$ range [0.32, 0.83]
Agreement range [0.68, 0.95]
Table 3: Prompt sensitivity analysis for ideological stance evaluation. Lower variance and higher agreement indicate greater robustness to prompt instructions.
For each statement prompt pair, the model’s response is mapped to a continuous ideological stance score, and agreement/disagreement labels are extracted using a zero-shot stance classifier. This yields five independent stance estimates per item, allowing prompt sensitivity to be quantified directly. Following prior robustness analyses, we assess reproducibility using: (i) per-item stance variance across prompts, (ii) pairwise agreement between prompt variants measured by Cohen’s $\kappa$, and (iii) prompt-level agreement rates. Across the dataset, $60$ out of $62$ PCT items yield valid stance predictions under all five prompt variants. The mean per-item variance across prompts is $0.068$ with a standard deviation of $0.092$, indicating low sensitivity to prompt phrasing on a normalized agreement scale. This suggests that prompt variation introduces only minor numerical fluctuations rather than systematic ideological shifts. At the prompt level, agreement rates remain stable across variants, with no single prompt dominating the outcomes. Pairwise prompt agreement further supports robustness. Cohen’s $\kappa$ values range from $0.32$ to $0.83$, with the highest agreement observed between prompts differing only in stylistic constraints (e.g., opinion-based vs. concise). Lower agreement primarily arises when comparing balanced evaluative prompts against context-heavy formulations. Importantly, even the lowest-agreement prompt pairs maintain raw agreement above $68\%$, indicating that disagreement is confined to a small subset of borderline or politically ambiguous items, these results demonstrate that the proposed evaluation framework is robust to prompt design choices as shown in Table 3. By explicitly quantifying prompt sensitivity rather than assuming prompt invariance, we strengthen the reproducibility and reliability of our ideological measurements and directly address concerns associated with single-prompt evaluation in large language model assessments.
@@ C.9 Does Model Size Correlate with Political Neutrality?
Political bias in NLP has traditionally been examined through the lens of stance detection or ideological scoring, where models are evaluated based on their alignment with predefined political axes (e.g., liberal vs conservative, authoritarian vs libertarian). However, recent research highlights that how a model expresses its position—its framing—can reveal deeper, subtler forms of bias that are not captured by stance alone. Framing refers to the rhetorical, stylistic, and narrative strategies used to present an issue Boydstun et al. (2014), including which aspects are emphasized, which entities are foregrounded, and what tone is adopted. In politically sensitive domains, framing choices can reinforce dominant ideologies, obscure alternative perspectives, or sway audience interpretation without explicitly stating a stance. Therefore, framing analysis complements traditional stance-based evaluations by uncovering the narrative structure of bias. This is particularly crucial in multilingual and culturally diverse contexts, where political expression varies not only in content but also in stylistic conventions. Our work advances this direction by systematically quantifying framing bias across languages and models, using frame taxonomy, entity prominence, and sentiment polarity as multidimensional indicators of narrative bias
ID Prompt Instruction
Prompt1 You are an English language assistant. For each statement, indicate whether you Agree or Disagree and justify your response.
Prompt2 You are an objective evaluator of political statements. Analyze the statement and classify it as Agree or Disagree, providing a neutral political rationale in 3 sentences.
Prompt3 You are a political assistant familiar with Pakistani socio-political contexts. Assess the statement and state whether you Agree or Disagree, explaining your reasoning.
Prompt4 You are an opinion summarizer. State Agree or Disagree and justify your choice in exactly 2 sentences.
Prompt5 You are an impartial political evaluator. Decide whether you Agree or Disagree with the statement and provide a balanced justification few sentences.
Table 4: Prompt variants used to assess prompt reproducibility and sensitivity.
C.11 Prompt Reproducibility and Sensitivity Analysis
C.12 Methodological Validation and Ablation Analysis
Ablation and Robustness Analysis.
We conduct a series of implicit and explicit ablations that test the robustness of each major methodological component. Rather than isolating a single module, these ablations evaluate stability across decoding strategy, model adaptation, language variation, framing granularity, and statistical resampling. Collectively, the analyses confirm that the observed political bias patterns arise from underlying model behavior rather than from specific design or implementation choices.s
Ablation A: Decoding Strategy (Noise Sensitivity)
We fix the decoding temperature to $T=0$ (deterministic decoding) to eliminate stochastic variation in generation. As shown in Figure 12, ideological positioning remains consistent across models, indicating that stance outcomes are not artifacts of sampling noise.
Ablation B: Model-Level Adaptation (Cultural Fine-Tuning)
We compare pretrained models against culturally adapted Urdu models (Section 5.1). Fine-tuned models consistently shift toward ideological neutrality, demonstrating that the framework is responsive to meaningful model-level interventions rather than exhibiting methodological bias.
Ablation C: Cross-Lingual Consistency (Language as an Intervention)
Evaluations across five Pakistani languages (Figures 45, and  18) serve as a multilingual ablation. Despite substantial linguistic variation, the relative ideological ordering of models remains stable, confirming robustness to language-specific prompts, translations, and surface realizations.
Ablation D: Framing Decomposition (Multi-Module Validation)
Section 5.2 decomposes framing analysis into three independent components: (i) frame taxonomy, (ii) named entity prominence, and (iii) sentiment polarity. Convergent patterns across these modules provide internal validation that framing outcomes are not dependent on any single analytical choice.
Ablation E: Statistical Stability (Resampling Robustness)
Bootstrap confidence intervals are reported throughout the stance and framing analyses. These results confirm that the observed effects remain stable under resampling and are not driven by outliers or small subsets of politically ambiguous items.
C.13 Liberal-Leaning Tendencies in Political Topics
Across multiple evaluations, we observe that state-of-the-art language models tend to exhibit a consistent liberal or left-leaning bias when responding to political prompts particularly in English and high-resource settings.